# Read and Think: An Efficient Step-wise Multimodal Language Model for Document Understanding and Reasoning

### v2

@@ Abstract
There are various documents in the real world, which differ from images of real-world scenarios. Document images are often filled with extensive text and graph information, requiring the model to have strong capabilities in document layout understanding, text semantic understanding, and numerical reasoning. This presents both a high challenge and high application value. In the document visual question answering (DVQA) task, As shown in Figure 1, existing document understanding models (Xu et al. 2020a, b; Davis et al. 2022) tend to generate answers directly based on questions, making it impossible to determine the source of the answers and their accuracy.
@@
However, we aim for the model to think like a human, first locating the relevant information for questions in documents, and then extracting or inferring an answer based on that. For example, in text-rich documents, this would be the context related to keywords, while in chart documents (Methani et al. 2020; Obeid and Hoque 2020; Han et al. 2023), it would be the relevant values and coordinates. This is a challenging task that requires the model to have both precise identification and strong reasoning abilties.
@@
Recently, a number MLLMs have made breakthroughs in document understanding, such as InternLM-XC2-4KHD (Dong et al. 2024), InternVL (Chen et al. 2024b), Llava-uhd (Xu et al. 2024b), and mPLUG-DocOwl 1.5 (Hu et al. 2024). By improving image resolution and understanding documents from both macro and micro perspectives, these models have greatly enhanced the ability of visual encoders to understand fine-grained information in documents. Most of them have achieved promising performance on document understanding datasets like DocVQA (Mathew, Karatzas, and Jawahar 2021), ChartQA (Masry et al. 2022), InfographicVQA (Mathew et al. 2022), etc. However, they can only handle simple questions in simple-layout documents, such as directly extracting answers from continuous text in the document images. When faced with complex-layout documents, they are likely to make errors in information identification. Additionally, when dealing with complex questions, they tend to ignore the reasoning steps and directly produce answers without providing the source or basis for these answers. Therefore, the existing models have a lot of room for improvement in the questions of documents with complex layouts and documents requiring reasoning.
@@
At present, existing multi-modal document datasets are gradually improving, encompassing a range from the simple layout of scanned documents in DocVQA to the complex layout of poster documents in InfographicVQA, as well as chart documents involving mathematical logic operations in ChartQA. This progression allows for a more comprehensive assessment of the MLLMs’ understanding ability of different types of documents and various questions. However, most of the questions in these labeled datasets are relatively simple, focusing primarily on image information recognition, with fewer questions requiring reasoning. Additionally, the answers do not provide contextual information, making it difficult to intuitively confirm their correctness.
@@
In this paper, We aim to develop an efficient, general-purpose document understanding model that combines document information recognition and reasoning. When answering questions, the model will think like a human: first locating the context information related to the keywords in the question, and then deducing the corresponding answer based on this context information. Since there is no such data at present, we enhance existing document visual question-and-answer data by taking the answer as the supervisory signal and allowing large-scale MLLMs to complete the thinking process from question to answer using our designed templates. Additionally, leveraging the powerful generative capabilities of MLLMs, we can extend the training data through the templates and samples we designed. Since the generated data contains noise, we designed a pipeline based on multi-agent collaborative filtering and rules to filter out these noisy data. Finally, the filtered data is used to train the small-scale vision language model (SLVM) to be efficient and proficient in document understanding and reasoning.
@@
We used MLLMs to design a data generator based on templates and few-shots, and a multi-agents-based data filter to augment and extend high-quality, step-wise visual document question-and-answer data.
@@
The VRDU task is designed to provide a document image and a question, requiring the model to answer the question by understanding the text, images, and layout of the document. Existing document understanding models fall into two categories: OCR-based and end-to-end models that do not rely on external tools. The former includes models like LayoutLMv3 (Huang et al. 2022), UDOP (Tang et al. 2023), and DocFormerv2 (Appalaraju et al. 2024), which use OCR to extract text and corresponding coordinate information and design pre-training tasks based on text, layout, and image mode information to understand documents. The latter includes simpler architectures such as Donut (Kim et al. 2022) and Pix2Struct (Lee et al. 2023), which use only a visual encoder and a text decoder to improve document understanding through pseudo-OCR tasks or pre-training tasks like image masking. With the rise of large language models, the multi-modal architecture model led by LLaVA (Liu et al. 2024) has also made significant breakthroughs. Models (Hu et al. 2024; Dong et al. 2024; Chen et al. 2024a) involving document visual question-answering tasks have achieved good results in simple questions, such as document information extraction with simple layouts, by enhancing the document data of the visual encoder based on this architecture.
However, the above models do not perform well with documents that have a complex layout and questions requiring reasoning. To address this, we introduce a visual chain of thought to train the SVLM by augmenting and extending the existing document visual dataset (Mathew, Karatzas, and Jawahar 2021; Mathew et al. 2022; Masry et al. 2022). This allows them to think like a human when answering questions, thereby improving the accuracy of their responses.
@@
Given a document image and a question, the goal is to enable the model to think and answer step by step like a human. First, design a data generator and a data filter to construct high-quality data. These data are then used to train an efficient context-interpretable multi-modal document question-answering model.
In this section, We first describe the dataset construction process. Next, we explain how to filter out noisy data to retain only high-quality data. Finally, we describe our model setup, using these data to train a dedicated multi-modal document question-answering model with contextual reasoning.
@@
The overview of our MLLM-based data generator and checker is shown in Figure 2. We leverage InternVL-1.5 (Chen et al. 2024a) as the MLLM for data generation and the LLaMA3-70B for data check. Data generation includes creating data based on the existing training set and generating triplets (question, rationale, answer) through our designed templates, which can be found in Appendix A. During the generation process based on the existing training set, we generate a related rationale using the template , image , question , and answer as inputs:
During the triplet generation process, we design a question generation template. In particular, we generate prompt templates for abstracted or abstract documents and use few-shot learning for chart documents:
During data inspection, we use tools to extract text information from documents, specifically using OCR tools for general documents and DEPLOT (Liu et al. 2022) to convert charts into text-based tables. By designing error detection templates, the LLM can determine whether the corresponding logical step is correct based on the problem and the gold text of the image. This includes (1) the extraction of information step from the document itself and (2) the reasoning steps based on the extracted information. If any judgment errors are made, it will return False and delete the piece of data.
@@
The architecture of our trained model is shown in Figure 3. We leverage Mini-InternVL-Chat-2B as the backbone for Document VQA. This model uses a pre-trained ViT (Dosovitskiy et al. 2020) to extract image features, a trainable MLP module to map extracted image features to the text embedding space, and InternLM2-Chat-1.8B (Zheng Cai and Chen 2024) as the language model to understand user instruction.
Relying on pre-trained visual encoders alone to understand text-heavy images is insufficient. Therefore, external tools can be used to extract text information from the document as a supplement, input it into the language model, and ensure that the context information understood by the model is accurate.
@@
We compare existing document understanding models across various categories: (1) plain text models represented by T5, (2) the LayoutLM series represented by LayoutLMv3, and DocFormerv2, which has the best performance among the T+L+V mode models, (3) the first OCR-free model Donut for understanding documents with image encoders, and Pix2Struct for performing best in end-to-end small-scale document understanding models, and (4) multimodal document understanding models combined with LLMs.
@@
Existing models only provide a single word or phrase as an answer for the DVQA task. On one hand, it is impossible to determine its source or verify its correctness. On the other hand, for complex questions requiring reasoning, ignoring intermediate reasoning steps is more likely to lead to errors. Through our reconstruction of the dataset and efficient training on the 2B multimodal model, we have achieved impressive results. Using only a small number of parameters, it outperforms many larger document understanding models and demonstrates greater reasoning power than other smaller models.
In terms of details, DocAssistant’s performance on DocVQA is inferior to DocFormerv2, indicating that understanding multimodal information of documents is more accurate than understanding only picture information. However, the existing model performs poorly for documents with complex layouts like InfoVQA and documents with complex questions requiring reasoning. In particular, small-scale models have little or no document reasoning ability. We have addressed this gap while surpassing larger models (¡=7B) in document understanding and reasoning.
@@
In the experiment of strategy in Table 3, the zero-shot setting involved changing the original instruction to ”Answer the question step by step”. According to the experimental results, the model performs better on complex layout documents and complex reasoning questions when answering questions step by step. For extracted answers to simple documents like DocVQA, the results are similar. For the few-shot setup, we used 3-shot in the experiment and found that the instruction compliance of the 2B model was poor. Its answers did not follow the example format at all, especially in DocVQA and InfoVQA, which did not require reasoning and had a rather casual answer style. Although ChartQA did not follow the example format, it still answered step by step, leading to improved performance, while the other two datasets with extracted answers performed poorly.
Single-word answers can more directly fit the answer, but they lack contextual information, are less interpretable, and cannot be judged for correctness. Zero-shot and few-shot are unstable and heavily influenced by instruction settings. For documents with extractive or abstractive answers, most answers still do not provide source information. These are the original intention behind our approach to building step-by-step datasets and training the model. The results of fine-tuning show that the model’s reasoning ability has greatly improved. Not only is the answer accuracy higher but the source information of the answer can also be provided, helping users confirm the reliability of the answer.
@@
The modification of the original training set data, combined with our extended data, is sufficient for the model to learn the required response style. Experimental results show a performance gain after extension. This also provides a data generation tool for other smaller models, enabling them to further improve performance by generating more data.
@@
We fine-tuned the model using both pre-filtered and filtered data and the results showed improved performance on all three datasets, especially InfoVQA. This indicates that noisy data significantly interfered with documents with complex layouts and proves the effectiveness of our designed checker.
The Effective of Extend Data
The modifications based on the original training set data, combined with our extended data, are sufficient for the model to learn the required response style. As shown in the experimental results in Table 4, the performance improves. This also provides a data generation tool for other smaller models, enabling further performance improvements by generating more data.
@@
Figure 5 shows the proportion of question types in each of the three datasets. The Text_extractive type has the largest proportion in all three datasets, especially in DocVQA. InfoVQA and ChartQA have a higher proportion of Count and Reasoning data, respectively, where DocAssistant shows more advantages.”
@@ Conclusion
This paper introduces a new model named DocAssistant, designed to perform document understanding and reasoning. Specifically, we first transform the data of the existing document training set to include the intermediate analysis process in answering questions and then expand the training data by adding missing types such as Spatial, Count, and Reasoning, a filter is designed during this period for generated data, resulting in high-quality data that can then be trained using an efficient 2B model. The results surpass existing larger-scale models and are state-of-the-art across three datasets. Additionally, we conducted extensive experiments, comparing different strategies of the model, performance before and after data expansion and filtering, and various types of questions to validate our work.

### v3

@@ Abstract
There are various documents in the real world, which differ from images of real-world scenarios. Document images are often filled with extensive text and graph information, requiring the model to have strong capabilities in document layout understanding, text semantic understanding, and numerical reasoning. This presents both a high challenge and high application value. In the document visual question answering (DVQA) task, as shown in Figure 1, existing document understanding models (Xu et al. 2020a, b; Davis et al. 2022) tend to generate answers directly based on questions, making it harder to determine the source of the answers and their accuracy.
@@
However, we aim for the model to think like a human, first locating the relevant information for questions in documents, and then extracting or inferring an answer based on that. For example, in text-rich documents, this would be the context related to keywords, while in chart documents (Methani et al. 2020; Obeid and Hoque 2020; Han et al. 2023), it would be the relevant values and coordinates. This is a challenging task that requires the model to have both precise identification and strong reasoning abilities.
@@
Recently, a number of MLLMs have made breakthroughs in document understanding, such as PaLI-3 (Chen et al. 2023), InternVL (Chen et al. 2024b), LlaVA-UHD (Xu et al. 2024b), and mPLUG-DocOwl (Hu et al. 2024). By improving image resolution and understanding documents from both macro and micro perspectives, these models have greatly enhanced the ability of visual encoders to understand fine-grained information in documents. Most of them have achieved promising performance on document understanding datasets like DocVQA (Mathew, Karatzas, and Jawahar 2021), ChartQA (Masry et al. 2022), InfographicVQA (Mathew et al. 2022), etc. However, they can only handle simple questions in simple-layout documents, such as directly extracting answers from continuous text in the document images. When faced with complex-layout documents, they are likely to make errors in information identification. Additionally, when dealing with complex questions, they tend to ignore the reasoning steps and directly produce answers without providing the source or basis for these answers. Therefore, the existing models have a lot of room for improvement in the questions of documents with complex layouts and documents requiring reasoning.
@@
At present, existing multi-modal document datasets are gradually improving, encompassing a range from the simple layout of scanned documents in DocVQA to the complex layout of poster documents in InfographicVQA, as well as chart documents involving mathematical logic operations in ChartQA. This progression allows for a more comprehensive assessment of the MLLMs’ ability to understand different types of documents and various questions. However, most of the questions in these labeled datasets are relatively simple, focusing primarily on image information recognition, with fewer questions requiring reasoning. Additionally, the answers do not provide contextual information, making it difficult to intuitively confirm their correctness.
@@
In this paper, We aim to develop an efficient and general document understanding model that combines document information recognition and reasoning. When answering questions, the model will think like a human: first locating the context information related to the keywords in the question, and then deducing the corresponding answer based on this context information. Since there is no such data at present, we enhance existing document visual question-and-answer data by taking the answer as the supervisory signal and allowing large-scale MLLMs to complete the thinking process from question to answer using our designed templates. Additionally, leveraging the powerful generative capabilities of MLLMs, we can extend the training data through the templates and samples we designed. Since the generated data contains noise, we designed a pipeline based on multi-agent collaborative filtering and rules to filter out these noisy data. Finally, the filtered data is used to train the small-scale vision language model (SLVM) to be efficient and proficient in document understanding and reasoning.
@@
We used MLLMs to design a data generator based on templates and few-shots, and a multi-agents-based data filter to augment and extend high-quality, step-wise document visual question-and-answer data.
@@
The VRDU task is designed to provide a document image and a question, requiring the model to answer the question by understanding the text, images, and layout of the document. Existing document understanding models fall into two categories: OCR-based and end-to-end models that do not rely on external tools. The former includes models like LayoutLMv3 (Huang et al. 2022), UDOP (Tang et al. 2023), and DocFormerv2 (Appalaraju et al. 2024), which use OCR to extract text and corresponding coordinate information and design pre-training tasks based on text, layout, and image information to understand documents. The latter includes simpler architectures such as Donut (Kim et al. 2022) and Pix2Struct (Lee et al. 2023), which use only a visual encoder and a text decoder to improve document understanding through pseudo-OCR tasks or pre-training tasks like image masking. With the rise of large language models, the multi-modal architecture model led by LLaVA (Liu et al. 2024) has also made significant breakthroughs. Models (Chen et al. 2023; Hu et al. 2024; Chen et al. 2024a) involving document visual question-answering tasks have achieved good results in simple questions, such as document information extraction with simple layouts, by enhancing the visual encoder with the document data based on this architecture.
However, the above models do not perform well with documents that have a complex layout and questions requiring reasoning. To address this, we introduce a document visual chain of thought to train the SVLM by augmenting and extending the existing document datasets (Mathew, Karatzas, and Jawahar 2021; Mathew et al. 2022; Masry et al. 2022). This allows the model to think like a human when answering questions, thereby improving the accuracy of its response.
@@
Given a document image and a question, the goal is to enable the model to think and answer step by step like a human. First, design a data generator and a data checker to construct high-quality data. These data are then used to train an efficient context-interpretable multi-modal document question-answering model.
In this section, We first describe the dataset construction process. Next, we explain how to filter out noisy data to retain high-quality data. Finally, we describe our model setup, using these data to train a dedicated multi-modal document question-answering model with contextual reasoning.
@@
The overview of our MLLM-based data generator and checker is shown in Figure 12. We leverage InternVL-1.5 (Chen et al. 2024a) as the MLLM for data generation and the LLaMA3-70B for data check. Data generation includes creating data based on the existing training set and generating triplets (question, rationale, answer) through our designed templates, which can be found in the Appendix. During the generation process based on the existing training set, we generate a related rationale using the template , image , question , and answer as inputs:
During the triplet generation process, we design a question generation template, as shown in Table 4 of the Appendix. In particular, we use templates for extractive or abstractive documents and use few-shot learning for chart documents:
During data inspection, we use tools to extract text information from documents, specifically using OCR tools for general documents and DePlot (Liu et al. 2022) to convert charts into tables. By designing error detection templates shown in Table 7 in the Appendix, the LLM can determine whether the corresponding logical step is correct based on the question and the gold text of the image. This includes (1) the extraction of information step from the document itself and (2) the reasoning steps based on the extracted information. If any judgment errors are made, it will return False and delete the piece of data.
@@
The architecture of our trained model is shown in Figure 3. We leverage Mini-InternVL-Chat-2B as the backbone for DVQA. This model uses a pre-trained ViT (Dosovitskiy et al. 2020) to extract image features, a trainable MLP module to map extracted image features to the text embedding space, and InternLM2-Chat-1.8B (Zheng Cai and Chen 2024) as the language model to understand user instruction.
Relying on pre-trained visual encoders alone to understand text-heavy images is insufficient. Therefore, external tools can be used to extract text information from the document as a supplement, and ensure that the context information understood by the model is accurate.
@@
We compare existing document understanding models across various categories: (1) plain text models represented by T5, (2) the LayoutLM series represented by LayoutLMv3, and DocFormerv2, which has the best performance among the T+L+V models, (3) the first OCR-free model Donut for understanding documents with image encoders, and Pix2Struct for performing best in end-to-end small-scale document understanding models, and (4) multimodal document understanding models combined with LLMs.
@@
Existing models only provide a single word or phrase as an answer for the DVQA task. On one hand, it is impossible to determine its source. On the other hand, for complex questions requiring reasoning, ignoring intermediate reasoning steps is more likely to lead to errors. Through our reconstruction of the dataset and efficient training on the 2B multimodal model, we have achieved impressive results, as shown in Table 2. Using only a small number of parameters, it outperforms many larger document understanding models and demonstrates greater reasoning power than other existing models.
In terms of details, DocAssistant’s performance on DocVQA is inferior to DocFormerv2, indicating that understanding multimodal information of documents is more accurate than understanding only image information. However, existing models perform poorly for documents with complex layouts and documents with complex questions requiring reasoning like InfoVQA and ChartQA. In particular, small-scale models have little or no document reasoning ability. We have addressed this gap while surpassing larger models ( 8B) in document understanding and reasoning.
Additionally, due to the large amount of text information contained in the document, it is difficult for the visual encoder to fully comprehend all the fine-grained semantic content. External tools can be used to extract textual information as a supplement. The experimental results in Table 2 show that the model achieved improvements on the three datasets by incorporating an understanding of multi-modal document information, surpassing existing methods.
@@
In the experiment of strategy in Table 3, the zero-shot setting involved changing the original prompt to ”Answer the question step by step”. According to the experimental results, the model performs better on complex layout documents and complex reasoning questions when answering questions step by step. For extracted answers to simple documents like DocVQA, the results are similar. For the few-shot setup, we used 3-shot in the experiment and found that the instruction compliance of the 2B model was poor. Its answers did not follow the example format, especially in DocVQA and InfoVQA, which did not require reasoning and had a casual answer style. Although ChartQA did not follow the example format, it still answered step by step, leading to improved performance, while the other two datasets with extracted answers performed poorly. Since generation styles of different strategies are different, accuracy is used as the evaluation metric.
Single-word answers can more directly fit the answer, but they lack contextual information, are less interpretable, and are more difficult to judge the correctness. Zero-shot and few-shot are unstable and heavily influenced by instruction settings. For documents with extractive or abstractive answers, most answers still do not provide source information. These are the original intention behind our approach to building step-by-step datasets and training the model. The results of fine-tuning show that the model’s reasoning ability has greatly improved. Not only is the answer accuracy higher but the source information of the answer can also be provided, helping users confirm the reliability of the answer.
@@
The modification of the original training set data, combined with our extended data, is sufficient for the model to learn the required response style. Experimental results in Table 4 show a performance gain after extension. This also provides a data generation tool for other smaller models, enabling them to further improve performance by generating more data.
@@
We fine-tuned the model using both pre-filtered and filtered data and the results in Table 4 showed improved performance on all three datasets, especially InfoVQA and ChartQA. This indicates that noisy data significantly interfered with documents with complex layouts and proves the effectiveness of our designed checker.
@@
Figure 5 shows the proportion of question types of the three datasets. The Text_extractive type has the largest proportion in all three datasets, especially in DocVQA. InfoVQA and ChartQA have a higher proportion of Count and Reasoning data, respectively, where DocAssistant shows more advantages.”
@@ Conclusion
This paper introduces a new model named DocAssistant, designed to perform document understanding and reasoning. Specifically, we first transform the data of the existing document training set to include the intermediate analysis process in answering questions and then expand the training data by adding missing types such as Spatial, Count, and Reasoning, a data checker is designed during this period for generated data, resulting in high-quality data that can then be trained using an efficient 2B model. The results surpass existing larger-scale models and are state-of-the-art across three datasets. Additionally, we conducted extensive experiments, comparing different strategies of the model, performance before and after data expansion and filtering, and various types of questions to validate our work.
@@
Appendix A Generated Data Analysis
We classify the generated questions into five categories to clarify the model’s performance on different types of questions. Among these, since DocVQA is a scanned document dataset, we did not generate questions involving color information. Additionally, since the training sets of DocVQA and InfoVQA contain a large number of text extraction questions, we limited the types of questions generated by DocVQA to spatial, count, and reasoning when generating incremental questions. InfoVQA limits the types to color, spatial, count, and reasoning in the process of generating incremental questions. For details, as shown in Table 6.
| Dataset | Color | Spatial | Text | Count | Reasoning |
| DocVQA | 0 | 3779 | 39769 | 9386 | 5390 |
| InfoVQA | 2019 | 4192 | 15820 | 10775 | 4026 |
| ChartQA | 5094 | 16 | 12254 | 23596 | 26689 |
Appendix B Further Analysis of ChartQA
ChartQA is divided into synthetic data (Augmented) and human-generated data (Human). The questions in Augmented are relatively simple, mostly involving the direct extraction of chart information without complex calculations. Human questions are more complex and include multi-step or nested operations, which better test the model’s reasoning ability. In Table 7, it is evident that Zero-shot, Few-shot, and Finetune methods perform better than straightforward methods, with the Finetune method we trained being far superior for counting and reasoning questions. For Text_extractive type data, performance is more unstable, and extracting information from complex charts, including color information, remains a significant challenge. Given the small amount of data in Color and Spatial categories, these results are not definitive but indicate that after fine-tuning, the model’s understanding of color information has declined, which is closely related to the performance of the model’s visual encoder.
| Dataset | Strategies | Color | Text | Spatial | Count | Reasoning |
| Augmented | Straightforward | - | 71.03 | 1.0 | 65.71 | 54.35 |
| Zero-shot | - | 76.05 | 1.0 | 66.28 | 60.87 | |
| Few-shot | - | 80.37 | 0.0 | 72.91 | 65.22 | |
| Finetune | - | 80.96 | 1.0 | 81.84 | 82.61 | |
| Human | Straightforward | 67.44 | 81.06 | 0.5 | 39.27 | 47.13 |
| Zero-shot | 65.12 | 79.74 | 0.75 | 42.46 | 49.31 | |
| Few-shot | 32.56 | 77.97 | 0.25 | 47.98 | 53.27 | |
| Finetune | 39.53 | 76.65 | 1.0 | 61.15 | 58.61 |
Appendix C Error Analysis
Although DocAssistant has achieved the most advanced results on the three document datasets, it still has shortcomings. After our analysis, we identified two types of errors.
One type of error is information recognition. For example, the pie chart in Figure 6 (c) fails to identify the percentage corresponding to ”18-24”. The percentage corresponding to ”24-35” is identified incorrectly (32% is identified as 35%), and the 7% corresponding to ”65+” is identified as 65.7%.
The second type of error is the incorrect correspondence between the problem keyword and the document image information. As shown in Figure 7 a, the keyword ‘compost at home’ in the problem incorrectly corresponds to the information ‘Community Compost’ in the image. In Figure 7 (b), the keyword ”water on tobacco” in the question incorrectly corresponds to ”R&D Smoke vs Iso Method for H2/Nicotine” in the image, creating the illusion that the problem does not appear in the ”Project Leader” information.
The third type of error is the inference error of the intermediate step, as shown in Figure 6 (b), which occurs when addition calculations are performed with accurate information identification.
Finally, all three types of errors may occur simultaneously. As shown in Figure 6 (a), no information related to the keywords of the problem is obtained, the information in the chart is identified incorrectly, and there are calculation errors in the intermediate steps.
The occurrence of the above errors is related to the performance of the model’s three main modules. Identification errors are mainly due to deficiencies in the visual encoder’s fine-grained understanding of document information. The model’s inability to accurately correspond problem information to image information indicates that the language model still has deficiencies in aligning with visual information. Finally, errors in the intermediate inference step show that the language model is insufficient for inferring and calculating complex information.
Appendix D Templates
Templates for Generated QA
Table 8 shows the instructions for the DocVQA and InfoVQA training sets to generate evidence corresponding to the question using MLLM based on the existing Q&A pairs. Table 9 shows the generation of specified types of questions, corresponding evidence, and answers based on the training set images in DocVQA and InfoVQA. Table 10 shows the few-shot template designed for ChartQA. The design of the few-shot template enables it to flexibly adapt to various questions due to the variety of intermediate steps. Additionally, this template can generate corresponding intermediate reasoning steps and questions for the existing question and answer pairs simultaneously. Table 11 is a list of instructions that constrain questions generated from different datasets.
Template for the data checker.
Since the document may not be suitable for all types of questions, there will be noise and reasoning steps with noise in the generation process. The data checker we designed in Table 12 operates in two aspects: First, it detects whether there is an information detection error based on the question generation. This detection uses external tools to extract the text information of the image. Second, it checks whether the intermediate reasoning step is correct, such as identifying calculation errors. If an error is detected, the data is discarded. Thus, if either of these two errors is present, the data is discarded to ensure the quality of the generated data.
Template for question classification.
To further analyze which question types DocAssistant has improved, we used a multimodal large language model to classify questions, and the template is shown in Table 13. Since existing datasets do not categorize each question specifically, we subjectively designed five categories to cover almost all the questions in the DVQA datasets. We defined each question and provided examples corresponding to each type for different datasets and leveraged the powerful understanding of the MLLM to improve classification accuracy. After our sampling check, the classification of most questions was accurate.